### v1

@@ 4 Practitioners’ Perspectives
In order to ground the concept of ABPM—as well as the “agents for BPM versus BPM for agents”-dichotomy—in real-world perspectives, we conducted a series of interviews with relevant BPM practitioners. Several process management software vendors, including UiPath, Salesforce, IBM, and Workday, use the term agentic AI to describe the interplay between genAI and agents. Thus, we leveraged the term during the interviews as a starting point to discuss the impact of agents on process management, expecting it to resonate effectively with the interviewees. Our study leverages a qualitative research design to examine the perspectives of professionals in business process management and automation, focusing on the integration and impact of agentic AI. It investigates the participants’ understanding, expectations, and concerns related to agentic AI’s autonomy, adaptability, human collaboration, and governance. Alongside discussing the characteristics of agentic AI, the study places greater emphasis on its impact on practitioners’ organizations. Therefore, the study also explores to what extent Agentic Business Process Management as a methodological framework is needed to address management needs that arise from greater software autonomy.
4.1 Interview Methodology
@@
To systematically examine the perceptions and attitudes towards agentic AI, a qualitative content analysis approach [30] was followed. Semi-structured interviews were conducted with professionals across various industries, focusing on their understanding, expectations, and concerns regarding autonomy, adaptability, human collaboration, and governance of agentic AI. The process began with transcription, familiarization with the interview data and analysis using a deductive-inductive coding approach. Initial deductive categories were derived from the research questions, while additional inductive subcategories emerged from the data, allowing themes to derive directly from the participants’ responses. A total of 22 participants from various industries and roles were interviewed (see Fig. 2). Sessions lasted 60 to 90 minutes, with some conducted in pairs or groups due to accessibility constraints.
@@
4.2 Results
@@
5 Discussion
The findings show that organizations need a clear management framework, which we define as Agentic Business Process Management, to successfully introduce agents into their business processes. This framework aims to help organizations adopt agents effectively while ensuring humans stay informed and involved. The interview results underscore the two main pillars of agentic BPM. First, it links business context, guardrails, and human-agent collaboration to the deployment and execution of agents, ensuring these technologies help achieve process goals.
@@ 5 Discussion
Organizations should define clear business goals for agents, specify the expected benefits, and estimate the costs associated with its introduction and maintenance.
@@
Second, it ties customization, risk management, and adoption, to the application of agent-based concepts for effectively designing and analyzing processes involving autonomous software agents.
@@ 6 Conclusion
The rise of generative AI has led researchers and software vendors to explore its integration with agents, creating the concept of agentic AI. However, the impact of this technology and its management in critical processes remains unclear. Therefore, this paper set out to explore how BPM for agents—a methodological concept we define as agentic BPM—can assist organizations in addressing these challenges. Interviews of BPM practitioners highlight that the emerging generation of software agents may help automate repetitive tasks and enable humans to focus on meaningful work but also underline risks like bias, over-reliance, job loss, and lack of transparency. This emphasizes the importance of developing a methodology that ensures ethical use of agents through strong governance and active human oversight. The initial qualitative study presented in this paper highlights the need for broader research. Future studies should include participants with varied backgrounds, monitor the actual adoption of new generations of software agents (such as LLM-based agents), explore early adopters’ use cases, and assess the value of a defined agentic BPM framework. It should also investigate the application of agents in specific industries and develop strategies for its responsible adoption, balancing expected performance benefits with requirements regarding accountability and long-term system maintainability.

### v2

@@ 3 Methodology
Research design. Our study leverages a qualitative research design to examine the perspectives of professionals in business process management and automation, focusing on the integration and impact of agentic AI. It investigates the participants’ understanding, expectations, and concerns related to agentic AI’s autonomy, adaptability, human collaboration, and governance. Alongside discussing the characteristics of agentic AI, the study places greater emphasis on its impact on practitioners’ organizations. Therefore, the study also explores to what extent agent governance is required to address management needs that arise from greater software autonomy.
@@
To systematically examine the perceptions and attitudes towards agentic AI, a qualitative content analysis approach [22] was followed. Semi-structured interviews were conducted with professionals across various industries, focusing on their understanding, expectations, and concerns regarding autonomy, adaptability, human collaboration, and governance of agentic AI. The process began with transcription, familiarization with the interview data and analysis using a deductive-inductive coding approach. Initial deductive categories were derived from the research questions, while additional inductive subcategories emerged from the data, allowing themes to derive directly from the participants’ responses.
@@
We began by reviewing all the interview data to understand the context and identify recurring themes. Next, we applied open coding by labeling segments of text with basic descriptive codes (e.g., benefits, risks, and governance challenges related to agentic AI). We then grouped similar codes into broader categories during axial coding (e.g. ”automation potential” and ”workflow optimization” to ”benefits”), revealing patterns and common themes across the interviews. Finally, we refined and integrated these themes into a coherent framework during selective coding, ensuring that key insights were clearly defined. A total of 22 participants with experience with automation and AI technologies from various industries and roles were interviewed by the first author. In Figure 2, an overview of the participants demographic characteristics is provided, including their professional roles and the industries and countries in which they are employed. Sessions lasted 60 to 90 minutes, with some conducted in pairs or groups due to accessibility constraints.
4 Results
@@
5 Recommendations for ABPM Adoption
The findings reveal that organizations require a clear and structured management framework to successfully integrate AI agents into their business processes while ensuring that human actors remain informed, involved, and empowered. Insights from the expert interviews led to a set of practical recommendations and actions that organizations should consider during implementation. These span six critical focus areas: business context, operational guardrails, human-agent collaboration, customization, risk management, and adoption.
@@ 5 Recommendations for ABPM Adoption
Organizations should define specific goals for the use of AI agents, clarify the expected benefits, and assess the short- and long-term costs associated with their deployment and maintenance.
@@
In addition, the implications for BPM need to be considered. To bridge traditional BPM (focused on structure, control, and standardization) with agentic AI (emphasizing autonomy, adaptability, and self-learning), following actions are proposed to ensure their effective coexistence and alignment.
Human-agent balance.
Organizations should strive to design workflows that strategically leverage both AI capabilities and human expertise. This balance is essential for optimizing efficiency while preserving human judgment and creativity.
Human involvement.
As intelligent agents become integral to business processes, organizations must clarify whether human actors will collaborate with agents as peers or take on supervisory roles with selective intervention.
Process structures.
To fully realize the potential of self-adaptive agents, organizations should consider evolving beyond static process models and enabling more flexible, dynamic structures that can adjust in real time based on agent-driven insights.
Performance metrics.
Effective integration of agentic AI requires clear metrics and performance indicators. Organizations should develop governance frameworks that track the impact of agents on process efficiency, decision quality, and broader organizational outcomes.
6 Limitations and Future Work
Several limitations encountered during this study must be acknowledged. First, the sample was relatively small and predominantly drawn from experts at selected multinational organizations, which may not fully represent the diversity of experiences across various industries or among smaller enterprises. Second, the concept of agentic AI remains in its formative stage, as participants employed varied terminology and frameworks, making it difficult to develop a standardized definition. Consequently, the limited adoption and practical experience with this technology constrain the interpretation of the interview results.
Against this background, future research should focus on refining and empirically validating the conceptual frameworks for agentic AI, with particular emphasis on its ethical, legal, and technical implications. Additionally, quantitative studies and broader analyses across different industries are needed to assess the wider impact and evolution of agentic AI within various organizational contexts, and to develop robust governance structures that ensure transparent and accountable decision-making. Finally, future research should also examine how the governance of agentic AI with its core characteristics of autonomy and adaptability should differ from the current frameworks used for established technologies such as Robotic Process Automation.
@@ 7 Conclusion
The rise of generative AI has sparked growing interest in its integration with software agents, giving rise to the concept of agentic AI. While this emerging technology holds promise, its impact on business processes and the requirements for its effective management remain insufficiently understood. This paper addresses this gap by identifying key recommendations for organizations seeking to implement AI agents responsibly.
Interviews with BPM practitioners reveal that next-generation software agents can enhance efficiency by automating routine tasks and allowing humans to focus on higher-value activities. At the same time, practitioners express concerns about risks such as bias, over-reliance, job displacement, and a lack of transparency. These concerns underscore the need for a clear governance framework to guide agent deployment.
To support informed, involved, and empowered human participation, the study outlines six critical areas for organizational focus: clarifying business objectives, setting legal and ethical guardrails, enabling effective human-agent collaboration, tailoring agent behavior, managing operational risks, and ensuring safe, seamless adoption.
Beyond implementation, the findings have broader implications for Business Process Management. Agentic AI introduces new dynamics—autonomy, adaptability, and self-learning—that challenge traditional BPM principles centered on structure, control, and standardization. To align these approaches, organizations should take strategic steps: design workflows that balance human and agent input, redefine human roles as collaborators or supervisors, evolve static process models to support adaptive agent behavior, and implement robust metrics and governance frameworks to monitor performance and outcomes.
Together, these insights offer a practical roadmap for the responsible integration of AI agents. They help organizations harness the potential of agentic AI while safeguarding process integrity, organizational control, and stakeholder trust.